# Early-Cycle Internal Impedance Enables ML-Based Battery Cycle Life Predictions Across Manufacturers

### v1

@@ Acknowledgements
Appendix A Model Architecture Summary
While we only present results from a simple elastic net model in the manuscript, it is worth exploring the performance of other models — both simple and complex — for the data introduced in this study. Here, we perform the same analysis as in Section 3.1 in the manuscript using the cross-validation strategy detailed in the same section. Fig. 5 shows parity plots for different architectures using (i) $\Delta$Q(V) features, (ii) DCIR-based features, and (iii) all features. The same cycle pairing as in the manuscript has been used.
(a) Random forest model using $\Delta$Q(V)101-1 features
(b) Random forest model using DCIR features from cycles 2, 102
(c) Random forest model using all features
(d) Ridge regression model using $\Delta$Q(V)101-1 features
(e) Ridge regression model using DCIR features from cycles 2, 102
(f) Ridge regression model using all features
(g) Lasso regression model using $\Delta$Q(V)101-1 features
(h) Lasso regression model using DCIR features from cycles 2, 102
(i) Lasso regression model using all features
(j) NuSVR model using $\Delta$Q(V)101-1 features
(k) NuSVR model using DCIR features from cycles 2, 102
(l) NuSVR model using all features in the study
Figure 5: A parity plot comparison for multiple architectures. The first column of figures uses the original six feature $\Delta$Q(V) set from Severson et al., the second column of figures uses the novel augmented DCIR $\Delta$Q(V) features developed in this study, and the third column of figures uses all features. Each row represents a new architecture.
Appendix B Model Validation
In addition to the ”Leave One Triplicate Out” cross validation strategy employed in the manuscript, we perform traditional 5 fold cross-validation for direct comparison. The following parity plots and table are equivalent analyses to that presented in the main text, but with sklearn’s standard KFold cross validator with 5 splits (80/20 training/testing split). Shown in the parity plots of Fig. 6 are the compiled test set predictions of all 5 test folds, and the reported MAE and R2 are the averaged values of all test folds.
(a) Model performance using $\Delta$Q(V)101-1 features with 5fold cross-validation.
(b) Model performance using DCIR $\Delta$Q(V)102-2 features with 5fold cross-validation.
(c) Model performance using full feature set with 5fold cross-validation.
Figure 6: A parity plot comparison for elastic net model performance using (a) the original six feature $\Delta$Q(V) set from Severson et al, (b) the novel augmented DCIR $\Delta$Q(V) feature set, and (c) the full feature set described in the manuscript to predict EOL85.
@@ Appendix B Model Validation
$\Delta$Q(V)101-1 $359\pm 433$ (0.06) $259\pm 333$ (-0.01) $134\pm 136$ (0.26)
Non-DCIR features $366\pm 438$ (0.03) $256\pm 331$ (0.01) $122\pm 113$ (0.44)
DCIR-based features $202\pm 168$ (0.80) $109\pm 119$ (0.85) $40\pm 38$ (0.94)
All features $163\pm 127$ (0.87) $145\pm 165$ (0.73) $46\pm 63$ (0.88)
Table 4: Model performance for Elastic Net architecture using a standard 5-fold cross validation technique
Appendix C Feature correlation
Feature correlation (on a log scale) for DCIR-based univariate features is shown in Fig 7. For simplicity, only data for one pulse has been shown. The Pearson correlation coefficient for each has been provided as well. Generally speaking, we do not observe high feature correlation with the end of life of a battery, suggesting that principal component analysis to reduce the feature space is a key component of the machine learning workflow.
Figure 7: Feature correlation (log scale) of univariate DCIR-based features used in the study. For simplicity, only features from the first DCIR pulse have been shown, computed from cycles 2 and 102. The six univaraite features are (a) minimum, (b) maximum, (c) mean, (d) variance, (e) skew, and (f) kurtosis of the mini $\Delta Q(V)$ curve curved obtained from the first DCIR pulse, as explained in the main manuscript. The Pearson’s correlation coefficient has been displayed on each plot as well.

### v2

@@ Acknowledgements
Appendix A Model Validation
In addition to the ”Leave One Triplicate Out” cross validation strategy employed in the manuscript, we perform traditional 5 fold cross-validation for direct comparison. The following parity plots and table are equivalent analyses to that presented in the main text, but with sklearn’s standard KFold cross validator with 5 splits (80/20 training/testing split). Shown in the parity plots of figs. 6(a)-6(e) are the compiled test set predictions of all 5 test folds. The reported MAE in table 4 are the averaged values of all test folds.
Figure 6: Parity plots showing the 5fold CV prediction accuracy of models trained on various feature sets: (a) the original six feature $\Delta$Q(V) set from Severson et al, (b) the original six feature $\Delta$Q(V) set and linear fit of the lifetime curve, (c) the novel augmented DCIR $\Delta$Q(V) feature set, (d) the augmented DCIR $\Delta$Q(V) and pulse resistances, and (e) the full combined feature set as described in the manuscript.
@@ Appendix A Model Validation
$\Delta Q(V)_{101-1}$ 359 259 134
Non-DCIR features 366 256 122
DCIR-based features 202 109 40
All features 163 145 46
Table 3: Model performance for Elastic Net architecture using a standard 5-fold cross validation technique.
Appendix B Feature selection
The feature space was dimensionally reduced using Sequential Feature Selection (SFS). For each cross-validation scheme considered in this work, Table 4 shows the number of features retained and that subsequently underwent Principle Component Analysis (PCA).
CV scheme Feature space dimension
Leave one triplicate out 20
Leave one operating condition out 12
Leave one manufacturer out 10
Table 4: Outcome of sequential feature selection showing final feature space dimensions for each cross-validation (CV) scheme.
Appendix C Data quality and end-of-life labeling
For the results presented in this work, end-of-life was defined as the number of cycles required to reach 85% State of Health (SOH). The following figures provide an overview of the individual lifetimes for each cell with the end-of-life point indicated by a vertical line at the cycle considered to be end-of-life. Note that a few of the cells (e.g., Sam50E at 25 degrees C, 3.3 - 4.2 V) did not fully reach 85%, and a linear fit of the last 100 cycles was used to extrapolate to 85% SOH. For these cells, the EOL label should be taken with caution and may be overestimating the actual EOL if the cell were to ”nosedive” and experience rapid loss in capacity at some point in the extrapolation window. That said, due to the known challenges with data scarcity, we believe linear extrapolation provides a reasonable estimation of the EOL for these cells and the benefits of the additional data outweigh possible risks of overestimation from the methodologies used.
Figure 7: Overview of individual cell lifetimes with labeled end-of-life denoted with vertical line.
Figure 8: Overview of individual cell lifetimes with labeled end-of-life denoted with vertical line.
Figure 9: Overview of individual cell lifetimes with labeled end-of-life denoted with vertical line.
Figure 10: Overview of individual cell lifetimes with labeled end-of-life denoted with vertical line.
Figure 11: Overview of individual cell lifetimes with labeled end-of-life denoted with vertical line.
Figure 12: Overview of individual cell lifetimes with labeled end-of-life denoted with vertical line.
Appendix D Feature correlation
Feature correlation (on a log scale) for all features considered in this work are shown in Figs 13 - 27. Generally speaking, we do not observe high feature correlation with the end of life of a battery, suggesting that principal component analysis to reduce the feature space is a key component of the machine learning workflow.
Figure 13: Feature correlation of the traditional Q(V) feature set.
Figure 14: Feature correlation of the p1p2 feature set.
Figure 15: Feature correlation of the DCIR Q(V) feature set for pulse0.
Figure 16: Feature correlation of the DCIR Q(V) feature set for pulse1.
Figure 17: Feature correlation of the DCIR Q(V) feature set for pulse2.
Figure 18: Feature correlation of the DCIR Q(V) feature set for pulse3.
Figure 19: Feature correlation of the DCIR Q(V) feature set for pulse4.
Figure 20: Feature correlation of the DCIR Q(V) feature set for pulse5.
Figure 21: Feature correlation of the DCIR Q(V) feature set for pulse6.
Figure 22: Feature correlation of the DCIR Q(V) feature set for pulse7.
Figure 23: Feature correlation of the DCIR Q(V) feature set for pulse8.
Figure 24: Feature correlation of the DCIR Q(V) feature set for pulse9.
Figure 25: Feature correlation of the DCIR resistances of cycle 102 and 2 for pulse0, pulse1, and pulse2.
Figure 26: Feature correlation of the DCIR resistances of cycle 102 and 2 for pulse3, pulse4, and pulse5.
Figure 27: Feature correlation of the DCIR resistances of cycle 102 and 2 for pulse6, pulse7, pulse8, and pulse9.